# MetaSym: A Symplectic Meta-learning Framework for Physical Intelligence

### v2

@@ 4.3 Quadrotor
For our final benchmark, we evaluate on a quadrotor, a canonical yet challenging system due to its floating-base dynamics and discontinuous orientation representations [40, 41]. The floating base induces internal covariate shifts during training, as phase-space magnitudes could vary across time. Regarding its orientation representation, Euler angles suffer from singularities, while quaternions impose a unit-norm constraint that is typically handled via loss penalties, often leading to local minima and unstable training. To address this, we represent the pose of the quadrotor (i.e., position and orientation) using the tangent space of the $\mathbb{SE}(3)$ group [42], which avoids the explicit enforcement of the unit norm constraint. The results in Figure 4 verify that our framework effectively handles these challenges. Full simulation details are in Appendix B.3.
4.4 Benchmarks
We benchmarked MetaSym against three baselines: an autoregressive Transformer [18], a physics-informed dissipative Hamiltonian neural network (DHNN) [11], and a naive Multi-Layer Perceptron (MLP). Both MetaSym and the Transformer perform autoregressive next-timestep prediction, whereas the DHNN models the system’s symplectic and dissipative gradients, relying on an integrator for trajectory unrolling. The MLP also predicts the next state directly. Unlike the others, the DHNN cannot predict quadrature measurements due to ill-defined derivatives involving complex Wiener terms [43], and its training proved less stable than that of MetaSym and the Transformer. All models were evaluated in an autoregressive rollout regime. As shown in Table 1, MetaSym consistently outperforms the alternatives in long-horizon prediction accuracy, particularly when using a 30-timestep context window. It achieves lower error accumulation and requires fewer parameters, which is especially advantageous for high-dimensional systems like spring-meshes. The variance present across all models is attributed to the noise that we artificially added to each network’s input in order to highlight MetaSym’s robustness and real-world applicability. In quantum dynamics, where real-time prediction and control are constrained by hardware latency and memory [44], MetaSym’s compact architecture offers a significant advantage over Transformer-based architectures [45].
@@ 4.4 Benchmarks
Spring Mesh System Quantum System Quadrotor
Models Param. Count Trajectory MSE (±$\pm\sigma$) Param. Count Trajectory MSE (±$\pm\sigma$) Param. Count Trajectory MSE (±$\pm\sigma$)
@@
QuantumLSTM [46] N/A N/A 244 0.891 (0.233) N/A N/A
@@ 5 Conclusion
In this work, we introduced MetaSym, a novel deep-learning framework that combines structure-preserving symplectic networks with an autoregressive, decoder equiped with meta-learning for modeling a wide range of physical systems. The core idea rests on striking a balance between strong physical priors—namely the intrinsic symplectic structure of Hamiltonian mechanics—and the flexibility required to capture non-conservative effects and heterogeneous system parameters. Our experimental results across diverse domains—ranging from high-dimensional spring-mesh dynamics to open quantum systems and robotics-inspired quadrotor models—demonstrated that MetaSym outperforms state-of-the-art baselines in both long-horizon accuracy and few-shot adaptation with smaller model sizes.
@@
From a geometric perspective, the SymplecticEncoder approximates canonical flows while preserving key invariants, significantly mitigating energy drift and ensuring robust long-term predictions. The encoder’s invertible design enforces time-reversal symmetry and reduces error accumulation. Meanwhile, the ActiveDecoder models departs from ideal Hamiltonian evolution through autoregressive prediction and meta-attention. The resulting architecture is computationally efficient, given that it does not require explicit numerical integration during inference, and, through meta-learning, it readily adapts to system variations with minimal additional data. This approach offers a scalable and unified framework for high-fidelity physics modeling in realistic settings with provable near-symplectic properties (see Appendix C.4).
@@
Limitations & Future Work: Considering MetaSym’s promising performance, we seek to investigate several future directions such as incorporating a fully symplectic network for modeling realistic physics by exploiting the underlying structure of dissipation and our control signals. MetaSym’s effectiveness on real world system data also remains to be investigated. Another natural extension of few-shot adaptation is online learning and control, since MetaSym can quickly adapt to new system configurations with minimal data. This could be leveraged in real-time control loops and model-based Reinforcement Learning (RL) algorithms. Finally, future research could focus on how to integrate the decoder’s adaptation step with RL or adaptive Model Predictive Control (MPC) frameworks, effectively enabling self-tuning controllers in rapidly changing environments.
Acknowledgements
The authors would like to thank Lucas Schorling, Vassil Atanassov, Federico Fedele and Vivek Wadhia for the useful discussions. P.V. is supported by the United States Army Research Office under Award No. W911NF-21-S-0009-2. A.P is supported by University of Oxford’s Clarendon Fund. M.T.M. is sup- ported by a Royal Society University Research Fellowship. N.A. acknowledges support from the European Research Council (grant agreement 948932) and the Royal Society (URF-R1-191150).
@@ Appendix A Meta Learning Setup
As mentioned in Section 3, we elaborate on the algorithmic pipeline of our SymplecticEncoder and ActiveDecoder modules. In particular, we outline the MAML-based meta-learning strategy for the SymplecticEncoder and the bi-level adaptation meta-learning for the ActiveDecoder. Figure 5 provides an interpretable insight into the effect of inner adaptation during the meta-update step and its effectiveness.
@@ A.1 Encoder
Following Section 3.1, Algorithm 1 provides a high-level overview of how the SymplecticEncoder is trained via meta-learning.
@@
0:  $\mathcal{D}=\{\mathcal{D}^{(1)},\dots,\mathcal{D}^{(N)}\}$: Training data from $N$ related systems. Each $\mathcal{D}^{(i)}$ is a trajectory of states $\bigl{(}\mathbf{q}_{t}^{(i)},\mathbf{p}_{t}^{(i)},\mathbf{u}_{t}^{(i)}\bigr{)}_{t=1}^{T}$.
1:  for epoch $=1\to N_{\mathrm{epochs}}$ do
2:     for each mini-batch of systems $B\subseteq\{1,\dots,N\}$ do
3:        optimizer_theta.zero_grad()
4:        inner_optimizer.zero_grad()
5:        for each system $i\in B$ do
6:           Split the trajectory $\mathcal{D}^{(i)}$ into:
7:            $\mathcal{I}_{\mathrm{adapt}}\subset\{1,\dots,\mathcal{T}^{(i)}_{\mathrm{adapt}}\},\quad\mathcal{I}_{\mathrm{meta}}=\{1,\dots,\mathcal{T}^{(i)}_{\mathrm{meta}}\}\setminus\mathcal{I}_{\mathrm{adapt}}$.
8:           $\mathcal{L}_{\mathrm{meta}}\leftarrow 0$
9:           Fast Adaptation: Adapt system-specific parameters
10:           $\theta_{SE}^{(i)}\leftarrow\theta_{SE}$.clone().detach() {Detach}
11:           for $k=1\to K$ do
12:              inner_optimizer.zero_grad()
13:              $\mathcal{L}_{\mathrm{inner}}^{(i)}\leftarrow 0$
14:              for $t\in\mathcal{I}_{\mathrm{adapt}}$ do
15:                 $\bigl{(}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{)}\;\leftarrow\;\Phi_{\theta_{\mathrm{SE}}}\!\Bigl{(}\mathbf{q}^{(i)}_{t},\,\mathbf{p}^{(i)}_{t},\,\mathbf{u}^{(i)}_{t};\,\theta_{SE}^{(i)}\Bigr{)}$
16:                 $\mathcal{L}_{\mathrm{inner}}^{(i)}\;+=\;\bigl{\|}\bigl{[}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{]}-[\mathbf{q}^{(i)}_{t+1},\mathbf{p}^{(i)}_{t+1}]\bigr{\|}^{2}$
17:              end for
18:              ($\mathcal{L}_{\mathrm{inner}}^{(i)}/\mathcal{T}_{adapt}^{(i)}).\texttt{backward()}$()
19:              inner_optimizer.step()  {Update $\theta_{SE}^{(i)}$ only}
20:           end for
21:           Meta Update: Update the global parameters $\theta_{AD}$
22:           for $t\in\mathcal{I}_{\mathrm{meta}}$ do
23:              $\theta_{SE}\leftarrow\theta_{SE}^{(i)^{*}}$.clone() {No Detach}
24:              $\bigl{(}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{)}\;\leftarrow\;\Phi_{\theta_{\mathrm{SE}}}\!\Bigl{(}\mathbf{q}^{(i)}_{t},\,\mathbf{p}^{(i)}_{t},\,\mathbf{u}^{(i)}_{t};\,\theta_{SE}\Bigr{)}$
25:              $\mathcal{L}_{\mathrm{meta}}\;+\!=\;\bigl{\|}\bigl{[}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{]}-[\mathbf{q}^{(i)}_{t+1},\mathbf{p}^{(i)}_{t+1}]\bigr{\|}^{2}$
26:           end for
@@
28:        ($\mathcal{L}_{\mathrm{meta}}/\mathcal{T}_{meta}^{(i)}).\texttt{backward()}$() {Eq. (3)}
29:        optimizer_theta.step()  {Update $\theta_{SE}$ only}
30:     end for
@@ A.2 Decoder
0:  $\mathcal{D}=\{\mathcal{D}^{(1)},\dots,\mathcal{D}^{(N)}\}$: Training data from $N$ related systems. Each $\mathcal{D}^{(i)}$ is a trajectory of states $\bigl{(}\mathbf{q}_{t}^{(i)},\mathbf{p}_{t}^{(i)},\mathbf{u}_{t}^{(i)}\bigr{)}_{t=1}^{T}$.
1:  for epoch $=1\to N_{\mathrm{epochs}}$ do
2:     for each mini-batch of systems $B\subseteq\{1,\dots,N\}$ do
3:        optimizer_theta.zero_grad()
4:        for each system $i\in B$ do
5:           Split the trajectory $\mathcal{D}^{(i)}$ into:
6:            $\mathcal{I}_{\mathrm{adapt}}\subset\{1,\dots,\mathcal{T}_{\mathrm{adapt}}^{(i)}\},\quad\mathcal{I}_{\mathrm{meta}}=\{1,\dots,\mathcal{T}_{\mathrm{meta}}^{(i)}\}\setminus\mathcal{I}_{\mathrm{adapt}}$.
7:           Inner loop: Adapt local parameters $\zeta_{i}$
8:           for $k=1\to K$ do
9:              for $t\in\mathcal{I}_{\mathrm{adapt}}$ do
10:                 inner_optimizer.zero_grad()
11:                 $\zeta_{i}$.randomize() {Reinitialize $\zeta_{i}$}
12:                 $\bigl{(}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{)}\;\leftarrow\;\Phi_{\theta_{\mathrm{AD}}}\!\Bigl{(}\mathbf{q}^{(i)}_{t:c},\,\mathbf{p}^{(i)}_{t:c},\,\mathbf{u}^{(i)}_{t:c};\,\theta_{AD},\,\zeta_{i}\Bigr{)}$
13:                 $\mathcal{L}_{\mathrm{inner}}^{(i)}\;\!=\;\bigl{\|}\bigl{[}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{]}-[\mathbf{q}^{(i)}_{t+1},\mathbf{p}^{(i)}_{t+1}]\bigr{\|}^{2}$
14:                 ($\mathcal{L}_{\mathrm{inner}}^{(i)}/\mathcal{T}_{adapt}^{(i)}).\texttt{backward()}$() {Eq. (4)}
15:                 inner_optimizer.step()  {Update $\zeta_{i}$ only}
16:              end for
@@
18:           Outer loop: Meta-update for the global parameters $\theta_{AD}$
19:           optimizer_theta.zero_grad() {No update to $\zeta_{i}$ now}
20:           $\mathcal{L}_{\mathrm{outer}}^{(i)}\leftarrow 0$
21:           for $t\in\mathcal{I}_{\mathrm{meta}}$ do
22:              $\bigl{(}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{)}\;\leftarrow\;\Phi_{\theta_{\mathrm{AD}}}\!\Bigl{(}\mathbf{q}^{(i)}_{t:c},\,\mathbf{p}^{(i)}_{t:c},\,\mathbf{u}^{(i)}_{t:c};\,\theta_{AD},\,\zeta_{i}^{*}\Bigr{)}$
23:              $\mathcal{L}_{\mathrm{outer}}^{(i)}\;+\!=\;\bigl{\|}\bigl{[}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{]}-[\mathbf{q}^{(i)}_{t+1},\mathbf{p}^{(i)}_{t+1}]\bigr{\|}^{2}$
24:           end for
25:           ($\mathcal{L}_{\mathrm{outer}}^{(i)}/\mathcal{T}_{meta}^{(i)}).\texttt{backward()}$() {Eq. (5)}
26:           optimizer_theta.step()  {Update $\theta_{AD}$ only}
27:        end for

### v3

@@ 4.3 Quadrotor
For our final benchmark, we evaluate on a quadrotor, a canonical yet challenging system due to its floating-base dynamics and discontinuous orientation representations (Sanyal & Roy, 2023; Zhou et al., 2019). The floating base induces internal covariate shifts during training, as phase-space magnitudes could vary across time. Regarding its orientation representation, Euler angles suffer from singularities, while quaternions impose a unit-norm constraint that is typically handled via loss penalties, often leading to local minima and unstable training. To address this, we represent the pose of the quadrotor (i.e., position and orientation) using the tangent space of the $\mathbb{SE}(3)$ group (Solà et al., 2021), which avoids the explicit enforcement of the unit norm constraint. MetaSym is pretrained using simulation data for a variety of different systems described in detail in in Appendix B.3. Importantly, the pretrained network is then fine-tuned via the query-value weight adaptation technique introduced with the ActiveDecoder to a real-world dataset (Mohajerin & Waslander, 2018). The results not only demonstrate our method’s efficacy and robustness on real-world systems but also the fast-adaptation capability of MetaSym that effectively diminishes the sim-to-real gap. Even though a detailed real-world deployment analysis is beyond the scope of the current paper, we performed an additional study to verify our model’s inference time. The results indicate that MetaSym can run on an NVIDIA RTX4090 in 1.551 ms for its full 30 time-step context-window.
@@
Spring-mesh System Quantum System Quadrotor
Models Param. Count Traj. MSE ($\pm\sigma$) Param. Count Traj. MSE ($\pm\sigma$) Param. Count Traj. MSE ($\pm\sigma$)
@@
QLSTM (Flurin et al., 2020) N/A N/A 244 0.891 (0.233) N/A N/A
Dissipative SymODEN 3.2M 32.054 (20.337) N/A N/A 3187 29.174 (9.547)
FNO 3.3M 25.187 (24.223) 33602 7.648 (0.411) 3336 27.932 (11.157)
iMODE 3.2M 23.858 (11.552) 202 7.667 (0.436) 3204 57.190 (20.903)
@@
4.4 Benchmarks
We benchmarked MetaSym against three baselines, an autoregressive Transformer (Vaswani, 2017), a physics-informed dissipative Hamiltonian neural network (DHNN) (Greydanus et al., 2019), a naive Multi-Layer Perceptron (MLP), the Dissipative SymODEN (Zhong et al., 2020), the Fourier neural operator (FNO) (Li et al., 2021), and iMODE (Li et al., 2023). Both MetaSym and the Transformer perform autoregressive next time-step prediction, whereas the DHNN models the system’s symplectic and dissipative gradients, relying on an integrator for trajectory unrolling. The MLP also predicts the next state directly. Unlike the others, the DHNN cannot predict quadrature measurements due to ill-defined derivatives involving complex Wiener terms (Milburn, 1987), and its training proved less stable than that of MetaSym and the Transformer. All models were evaluated in an autoregressive roll-out regime. As shown in Table 1, MetaSym consistently outperforms the alternatives in long-horizon prediction accuracy, particularly when using a $30$ time-step context window. It achieves lower error accumulation and requires fewer parameters, which is especially advantageous for high-dimensional systems like spring-meshes. The variance present across all models is attributed to the noise that we artificially added to each network’s input in order to highlight MetaSym’s robustness and real-world applicability. In quantum dynamics, where real-time prediction and control are constrained by hardware latency and memory (Reuer et al., 2023), MetaSym’s compact architecture offers a significant advantage over Transformer-based architectures (Vaidhyanathan et al., 2024).
@@ 5 Conclusion
In this work, we introduced MetaSym, a novel deep-learning framework that combines structure-preserving symplectic networks with an autoregressive, decoder equipped with meta-learning for modeling a wide-range of physical systems. The core idea rests on striking a balance between strong physical priors, namely the intrinsic symplectic structure of Hamiltonian mechanics, and the flexibility required to capture non-conservative effects and heterogeneous system parameters. Our experimental results across diverse domains, ranging from high-dimensional spring-mesh dynamics to open-quantum systems and real-world quadrotor systems, demonstrated that MetaSym outperforms SOTA baselines in both long-horizon accuracy and few-shot adaptation with smaller model sizes even for real-world systems.
@@
The SymplecticEncoder approximates canonical flows while preserving key invariants, significantly mitigating energy drift and ensuring robust long-term predictions. The encoder’s invertible design enforces time-reversal symmetry and reduces error accumulation. Meanwhile, the ActiveDecoder models departs from ideal Hamiltonian evolution through autoregressive prediction and meta-attention. The resulting architecture is computationally efficient, given that it does not require explicit numerical integration during inference, and through meta-learning, it readily adapts to system variations with minimal additional data. This approach offers a scalable and unified framework for high-fidelity physics modeling in realistic settings with provable near-symplectic properties (see Appendix C.4).
Additionally, while a substantial line of work has explored Meta-PINNs and parameter-aware operator learning to generalize across varying physical coefficients exist (Penwarden et al., 2023; Wang et al., 2021). These methods effectively adapt to new PDE instances, they typically enforce physical consistency via "soft" regularization terms in the loss function. This reliance on soft constraints often necessitates careful hyperparameter tuning of penalty weights and does not strictly guarantee the preservation of conservation laws during inference. In contrast, MetaSym enforces symplectic symmetries through "hard" architectural constraints within the encoder. This design ensures that the core conservative dynamics remain on the symplectic manifold by construction, independent of optimization stability, while relegating only the non-conservative dynamics to the adaptable decoder.
@@
Limitations & Future Work. It is important to note that MetaSym’s strongest guarantees (e.g. bounded energy drift, near-symplectic behavior, and long-horizon stability) rely on the availability of at least some representative conservative data for pretraining the SymplecticEncoder. When such data are unavailable, MetaSym can be interpreted as a hybrid physics-informed / data-driven model whose guarantees degrade continuously toward those of a fully black-box architecture. Considering MetaSym’s promising performance, we seek to investigate several future directions such as incorporating a fully symplectic network for modeling realistic physics by exploiting the underlying structure of dissipation and our control signals. MetaSym’s effectiveness on real-world system data also remains to be investigated. Another natural extension of few-shot adaptation is online learning and control, since MetaSym can quickly adapt to new system configurations with minimal data. This could be leveraged in real-time control loops and model-based Reinforcement Learning (RL) algorithms. Finally, future research could focus on how to integrate the decoder’s adaptation step with RL or adaptive Model Predictive Control (MPC) frameworks, effectively enabling self-tuning controllers in rapidly-changing environments.
@@ Appendix A Meta Learning Setup
As mentioned in Section 3, we elaborate on the algorithmic pipeline of our SymplecticEncoder and ActiveDecoder modules. In particular, we outline the MAML-based meta-learning strategy for the SymplecticEncoder and the bi-level adaptation meta-learning for the ActiveDecoder. Fig. 8 provides an interpretable insight into the effect of inner adaptation during the meta-update step and its effectiveness.
@@ A.1 Encoder
Following Section 3.1, Algorithm 1 provides a high-level overview of how the SymplecticEncoder is trained via MAML meta-learning. More specifically, the Model-Agnostic Meta-Learning (MAML) is a meta-learning framework designed to train models that can adapt quickly to new tasks using only a small amount of data. The core idea is to learn an set of parameters that is not task-specific, but task agnostic. As a model-agnostic method, it can be applied to any model trained with gradient descent, including neural networks for classification, regression (e.g., our case with MetaSym), reinforcement learning, and beyond. During meta-training, MAML simulates adaptation by sampling systems, performing inner-loop updates on each system, and then optimizing the initial parameters via meta-gradients computed across systems. This results in a model that learns how to learn, making it highly effective in few-shot scenarios where fast generalization from limited supervision is required.
The task-specific gradient steps are performed on parameters that are initialized from a shared-parameter set, denoted as $\theta_{SE}$, for each system. These inner-loop (fast adaptation) updates are applied to independent copies of the initial parameters, denoted $\theta_{SE}^{(i)}$, one for each task (i.e., each system $i$). Crucially, the original $\theta_{SE}$ must remain unchanged during these task-specific updates, as it serves as the point from which all adaptations are made.
Only after the fast adaptation steps across tasks are completed do we compute the meta-gradient, based on the post-adaptation performance on each task, and use it to update the original initialization $\theta_{SE}$ via the outer-loop (meta) update. This structure ensures that $\theta_{SE}$ is optimized for adaptability, enabling efficient transfer to unseen systems.
@@
[…]
2:for epoch =1=1 to $N_{\mathrm{epochs}}$ do
3:  for each mini-batch of systems $B\subseteq\{1,\dots,N\}$ do
4:   optimizer_theta.zero_grad()
5:   inner_optimizer.zero_grad()
6:   for each system $i\in B$ do
7:     Split the trajectory $\mathcal{D}^{(i)}$ into:
8:      $\mathcal{I}_{\mathrm{adapt}}\subset\{1,\dots,\mathcal{T}^{(i)}_{\mathrm{adapt}}\}$, $\mathcal{I}_{\mathrm{meta}}=\{1,\dots,\mathcal{T}^{(i)}_{\mathrm{meta}}\}\setminus\mathcal{I}_{\mathrm{adapt}}$.
9:     $\mathcal{L}_{\mathrm{meta}}\leftarrow 0$
10:     Fast Adaptation: Adapt system-specific parameters
11:     $\theta_{SE}^{(i)}\leftarrow\theta_{SE}\texttt{.clone().detach()}$ $\triangleright$ Detach
12:     for $k=1$ to $K$ do
13:      inner_optimizer.zero_grad()
14:      $\mathcal{L}_{\mathrm{inner}}^{(i)}\leftarrow 0$
15:      for $t\in\mathcal{I}_{\mathrm{adapt}}$ do
16:        $\bigl(\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr)\leftarrow\Phi_{\theta_{\mathrm{SE}}}\!\Bigl(\mathbf{q}^{(i)}_{t},\,\mathbf{p}^{(i)}_{t},\,\mathbf{u}^{(i)}_{t};\,\theta_{SE}^{(i)}\Bigr)$
17:        $\mathcal{L}_{\mathrm{inner}}^{(i)}\leftarrow\mathcal{L}_{\mathrm{inner}}^{(i)}+\bigl\|\,[\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}]-[\mathbf{q}^{(i)}_{t+1},\mathbf{p}^{(i)}_{t+1}]\bigr\|^{2}$
18:      end for
19:      $(\mathcal{L}_{\mathrm{inner}}^{(i)}/\mathcal{T}_{\mathrm{adapt}}^{(i)})\texttt{.backward()}$
20:      inner_optimizer.step() $\triangleright$ Update $\theta_{SE}^{(i)}$ only
21:     end for
22:     Meta Update: Update the global parameters $\theta_{SE}$
23:     for $t\in\mathcal{I}_{\mathrm{meta}}$ do
24:      $\theta_{SE}\leftarrow\theta_{SE}^{(i)*}\texttt{.clone()}$ $\triangleright$ No detach
25:      $\bigl(\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr)\leftarrow\Phi_{\theta_{\mathrm{SE}}}\!\Bigl(\mathbf{q}^{(i)}_{t},\,\mathbf{p}^{(i)}_{t},\,\mathbf{u}^{(i)}_{t};\,\theta_{SE}\Bigr)$
26:      $\mathcal{L}_{\mathrm{meta}}\leftarrow\mathcal{L}_{\mathrm{meta}}+\bigl\|\,[\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}]-[\mathbf{q}^{(i)}_{t+1},\mathbf{p}^{(i)}_{t+1}]\bigr\|^{2}$
@@
28:   end for
29:   $(\mathcal{L}_{\mathrm{meta}}/\mathcal{T}_{\mathrm{meta}}^{(i)})\texttt{.backward()}$ $\triangleright$ Eq. 4
30:   optimizer_theta.step() $\triangleright$ Update $\theta_{SE}$ only
@@ A.2 Decoder
[…]
2:for epoch =1=1 to $N_{\mathrm{epochs}}$ do
3:  for each mini-batch of systems $B\subseteq\{1,\dots,N\}$ do
4:   optimizer_theta.zero_grad()
5:   for each system $i\in B$ do
6:     Split the trajectory $\mathcal{D}^{(i)}$ into:
7:      $\mathcal{I}_{\mathrm{adapt}}\subset\{1,\dots,\mathcal{T}_{\mathrm{adapt}}^{(i)}\},\quad\mathcal{I}_{\mathrm{meta}}=\{1,\dots,\mathcal{T}_{\mathrm{meta}}^{(i)}\}\setminus\mathcal{I}_{\mathrm{adapt}}$.
8:     Inner loop: Adapt local parameters $\zeta_{i}$
9:     for $k=1$ to $K$ do
10:      for $t\in\mathcal{I}_{\mathrm{adapt}}$ do
11:        inner_optimizer.zero_grad()
12:        $\zeta_{i}.\texttt{randomize()}$ $\triangleright$ Reinitialize $\zeta_{i}$
13:        $\bigl(\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr)\leftarrow\Phi_{\theta_{\mathrm{AD}}}\!\Bigl(\mathbf{q}^{(i)}_{t:c},\,\mathbf{p}^{(i)}_{t:c},\,\mathbf{u}^{(i)}_{t:c};\,\theta_{AD},\,\zeta_{i}\Bigr)$
14:        $\mathcal{L}_{\mathrm{inner}}^{(i)}\leftarrow\bigl\|\,[\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}]-[\mathbf{q}^{(i)}_{t+1},\mathbf{p}^{(i)}_{t+1}]\bigr\|^{2}$
15:        $\bigl(\mathcal{L}_{\mathrm{inner}}^{(i)}/\mathcal{T}_{\mathrm{adapt}}^{(i)}\bigr)\texttt{.backward()}$ $\triangleright$ Eq. 5
16:        inner_optimizer.step() $\triangleright$ Update $\zeta_{i}$ only
@@
18:     end for
19:     Outer loop: Meta-update for the global parameters $\theta_{AD}$
20:     optimizer_theta.zero_grad() $\triangleright$ No update to $\zeta_{i}$ now
21:     $\mathcal{L}_{\mathrm{outer}}^{(i)}\leftarrow 0$
22:     for $t\in\mathcal{I}_{\mathrm{meta}}$ do
23:      $\bigl(\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr)\leftarrow\Phi_{\theta_{\mathrm{AD}}}\!\Bigl(\mathbf{q}^{(i)}_{t:c},\,\mathbf{p}^{(i)}_{t:c},\,\mathbf{u}^{(i)}_{t:c};\,\theta_{AD},\,\zeta_{i}^{*}\Bigr)$
24:      $\mathcal{L}_{\mathrm{outer}}^{(i)}\leftarrow\mathcal{L}_{\mathrm{outer}}^{(i)}+\bigl\|\,[\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}]-[\mathbf{q}^{(i)}_{t+1},\mathbf{p}^{(i)}_{t+1}]\bigr\|^{2}$
25:     end for
26:     $\bigl(\mathcal{L}_{\mathrm{outer}}^{(i)}/\mathcal{T}_{\mathrm{meta}}^{(i)}\bigr)\texttt{.backward()}$ $\triangleright$ Eq. 6
27:     optimizer_theta.step() $\triangleright$ Update $\theta_{AD}$ only